OpenStudy (prussianmaster):

Find all possible numbers, c, so that y={3x-2c if x is less than or equal to 0}{2x^2+x+5c^2 if x>0} is a continuous function

ganeshie8 (ganeshie8):

Hi

OpenStudy (prussianmaster):

Hello!

ganeshie8 (ganeshie8):

Have you tried anything yet ?

OpenStudy (prussianmaster):

What do you mean?

OpenStudy (prussianmaster):

I'm not sure how to even start the problem.

ganeshie8 (ganeshie8):

Me neither, let's read the problem again.

OpenStudy (prussianmaster):

I can draw the equation out

ganeshie8 (ganeshie8):

what does it mean for a function to be "continuous" ?

ganeshie8 (ganeshie8):

Is below function continuous ? 

OpenStudy (prussianmaster):

No.

ganeshie8 (ganeshie8):

why not ?

OpenStudy (prussianmaster):

The graph splits at a number, and it looks as if it has two values for the same place on the graph.

ganeshie8 (ganeshie8):

Right, basically we want the function to be smooth

ganeshie8 (ganeshie8):

It seems the expression for the function is changing at x = 0 ?

ganeshie8 (ganeshie8):

Before x = 0, the expression is 3x-2c after x = 0, the expression is 2x^2+x+5c^2

ganeshie8 (ganeshie8):

For the function to be continuous, we want both those expressions to be same at x = 0

ganeshie8 (ganeshie8):

plugin x = 0 and set them equal : 3(0) - 2c = 2(0)^2+0+5c^2

ganeshie8 (ganeshie8):

see if you can solve "c"

OpenStudy (prussianmaster):

Okay!

OpenStudy (prussianmaster):

I got 0 and -2/5.

OpenStudy (prussianmaster):

Awesome! Thank you!

ganeshie8 (ganeshie8):

yw!